Can you solve this problem 8. Find the perimeter and area of a right triangle if one leg measures 12 yd. and the other leg measures 35 yd. Students must show all work step by step to receive full credit for the problem. Include correct units with your solution.
Answer by rfer(16322)

A=(l*w)/2
A=(12*35)/2
A=210 sq yds
------------
a^2+b^2=c^2
12^2+35^2=144+1225=1369
sq rt 1369=37
------------
P=37+12+35
P=84 yds